# A Survey of Graph Retrieval-Augmented Generation for Customized Large Language Models

### v1

@@ III-A3 Knowledge Integration
The quality of retrieved content can drastically affect generation, with irrelevant or misleading data potentially having deleterious effects. LeanContext [97] addresses this issue by utilizing reinforcement learning to selectively choose sentences that are most pertinent to the query, effectively minimizing the context size and reducing computational costs. Similarly, SELF-RAG [17] introduces a self-reflection mechanism where the language model assesses its own generated and retrieved content, allowing for corrective measures to be implemented during the generation process. SKR [98] proposes a dynamic framework whereby LLMs can rely on their pre-trained knowledge for recognized queries, resorting to retrieval only when necessary. FILCO [99] enhances this by training a context filtering model that screens out irrelevant data, thereby mitigating hallucination risks. Recent research by Lyu et al. [100] introduces an evaluation method specifically designed to assess the importance of the retrieved content. Their findings suggest that by pruning or re-weighting parts of the retrieval corpus, RAG systems can enhance their performance without requiring additional training rounds. In scenarios where the retrieval corpus is limited, approaches such as Selfmem [101] construct a memory pool from LLM-generated results, using an iterative selection framework to enhance generative quality. SAIL [102] forms an instruction-tuning dataset based on retrieval outcomes, fine-tuning LLMs to ground responses in reliable content while excluding distracting elements.
Efficiency is another crucial concern, since encoding a large volume of retrieved passages is resource-intensive, leading to significant computation and memory overheads. To circumvent this, [103] adopts a pre-encoding strategy that stores dense representations of the text corpus. LUMEN [104] takes it a step further by conditioning the encoder on current inputs, and fine-tuning it specifically for this task to ensure high-quality pre-encoded memory representations. Innovative parallelism solutions, like PipeRAG [105], advance efficiency by enabling concurrent retrieval and generation, thus introducing alignment between retrieved content and real-time generation states. RAGCache [106] further optimizes resource usage by caching intermediate states within a GPU-resident knowledge tree, employing a prefix-aware replacement policy to maintain the most critical key-value tensors. To efficiently compress retrieval contents, Xu et al. [16] introduce an extractive compressor to identify useful sentences and an abstract compressor to distill summaries; this dual compression system refines textual summaries before they are integrated into the context. Similarly, TCRA-LLM [107] utilizes a token compression strategy, optimizing the semantic relevance of tokens to conserve computational resources.
@@ III-B1 Complex query understanding
Traditional RAG faces significant challenges in precisely answering complex queries, mainly due to the intrinsic limitation of its knowledge organization (vector database). Given a query, these RAG methods can only retrieve information from chunks containing anchor entities and are incapable of multi-hop reasoning. This limitation becomes more pronounced as granularity decreases, making it challenging to handle domain knowledge, as domain knowledge and complex logic often require multi-hop reasoning for effective understanding. Some traditional RAG methods have attempted to improve the complex query understanding, such as enhancing queries before retrieval [91], enhancing multiple query candidates for re-ranking [92], and using related metadata to provide richer information [85]. However, these methods are still limited by the chunking style of knowledge organization and are challenging to capture multi-hop information in complex queries effectively.
@@ III-C1 Definition of GraphRAG
GraphRAG can be formally defined as a subclass of RAG framework that leverage graph structure to organize and retrieve knowledge. Unlike traditional RAG methods, which rely on vector databases for knowledge organization, GraphRAG employs structural databases where graphs are used to model dependencies among knowledge pieces. This approach enhances the accuracy and efficiency of information retrieval, enabling more effective augmented generation for LLMs.
Specifically, GraphRAG systems can utilize graphs as either the carrier of knowledge or the indexing tools for efficient retrieval from chunked textual data. By modeling dependencies between nodes, GraphRAG enables the discovery of related knowledge pieces centered around a topic or anchor entity, ensuring comprehensive knowledge retrieval. Moreover, these connections support efficient search by navigating through relevant pathways and meanwhile pruning irrelevant information during the retrieval process.
@@ III-C2 Workflow of GraphRAG
  • Knowledge Organization The knowledge organization stage structures external information using graph-based methods, either through explicit knowledge representation (graphs as knowledge carriers) or indexing mechanisms (graphs for knowledge indexing). These approaches enable efficient, context-aware information retrieval.
  • Knowledge Retrieval. GraphRAG models employ graph-based planners (either learnable planners or graph algorithm-based planners) to retrieve relevant information based on the input query. These retrieval techniques not only consider the semantic similarity between the query and each text chunk, but also the logical coherence between the query type and the retrieved subgraph.
  • Knowledge Integration. Once the relevant knowledge is retrieved, the GraphRAG model integrates it with the input query to generate the final output. The goal of the integration process is to seamlessly incorporate the retrieved knowledge into the generated text, thereby enhancing its quality and informativeness. A key design consideration is how to preserve the richness of the retrieved subgraph information in the final text-based prompt without introducing redundancy or misplacing emphasis on less critical aspects of the textual descriptions.
@@ III-D Traditional RAG vs GraphRAG
GraphRAG provides several key advantages over traditional RAG systems, enhancing the capabilities of AI-driven information retrieval and generation. Below, we discuss these benefits and illustrate the comparison between traditional RAG and GraphRAG in Figure 4.
@@ III-D3 Efficiency and Scalability
GraphRAG systems built on fast graph databases can handle large-scale datasets efficiently. Graph databases are optimized for relationship-based queries, allowing for quick traversal of complex data structures. This efficiency translates to faster response times, especially for queries that require exploring multiple relationships. Research has shown that GraphRAG systems can generate LLM responses using 26% to 97% fewer tokens compared to traditional methods, indicating significant improvements in both speed and resource utilization. GraphRAG systems can more easily accommodate updates to the knowledge base. New information can be added as nodes or edges in the graph without requiring a complete reindexing of the entire knowledge base. This feature allows for real-time updates and ensures that the system always has access to the most current information. The graph structure allows GraphRAG to consider the context of a query more effectively. Instead of relying solely on keyword matching or vector similarity, it can leverage the semantic relationships in the graph to retrieve more relevant information. This context awareness leads to more accurate and pertinent responses to user queries.
@@ III-D4 Interpretability
The graph structure in GraphRAG allows for better visualization and understanding of how the system arrives at its responses. Users can trace the path of reasoning through the knowledge graph, seeing which entities and relationships were considered in formulating the answer. This transparency is crucial for building trust in AI systems and is especially valuable in fields like healthcare, finance, or legal applications where decision-making processes need to be auditable. By integrating LLMs with graph-based knowledge representation, GraphRAG can perform more sophisticated reasoning tasks. It can follow chains of logic across multiple nodes in the graph, combining information from various sources to answer complex queries. This capability allows for multi-hop question answering and the ability to draw insights that may not be explicitly stated in any single document.
@@
TABLE I: Comparison of three knowledge organization paradigms: $(i)$ vector databases in traditional RAG systems (“vector database”), $(ii)$ graphs for knowledge indexing (“index graphs”), and $(iii)$ graphs as knowledge carriers (“knowledge graphs”).
RAG Technique Traditional RAG GraphRAG
Knowledge Organization Paradigm Vector Database Index Graphs Knowledge Graphs
Advantages Explicit Reasoning on Graphs
Index Original Content
Logic-guided Retrieval
Community Detection
Prune Query-irrelevant Facts
Missing Fact Completion
Multi-hop Knowledge Edit
Disadvantages Sensitive to Chunk Granularity
Information Loss
@@ IV Knowledge Organization
A major distinction of GraphRAG techniques compared to traditional RAG methods lies in their ability to leverage graph structures for efficient knowledge organization, enhancing the effectiveness of query responses. In this setup, the LLM functions as an intelligent agent, while the graph structures support its ability to organize and integrate information more comprehensively and precisely. In real-world applications, external knowledge sources may include extensive text corpora, document collections, search results, historical user data, or interaction logs. Properly organizing these sources is essential, as direct retrieval can be prone to irrelevant results and missing contexts, leading to information overload, knowledge conflicts, and compromised comprehension.
@@ IV-A Graphs for Knowledge Indexing
The inherent challenges associated with the construction and maintenance of these graphs raises several research questions, giving rise to several research questions that shape current research directions: (1) Conciseness and Relevance: Ensuring the constructed graph captures only relevant relationships without overloading it with unnecessary connections is a significant challenge, thus facilitating effective recalling of relevant text chunks without redundancies, (2) Consistency and Conflict Resolution: Different data chunks may introduce conflicting information. It is crucial to resolve these conflicts and ensure that the graph remains consistent, reliable, and well-structured. Researchers are actively exploring tailored index graph structures to address these challenges effectively.
@@ IV-B Graphs as Knowledge Carriers
The paradigm of representing knowledge explicitly through graph structures is increasingly recognized for its effectiveness, offering several notable advantages: 1) Efficient retrieval of query-related knowledge facts. The explicit structure of knowledge graphs facilitates logic-guided chain retrieval, efficiently identifying missing facts while pruning the search space through reasoning paths; 2) Coherent multi-step reasoning over long spans. Previous studies have revealed the importance of the reasoning steps, i.e., lengthening the reasoning steps in prompts brought significant reasoning performance enhancement [108]. Knowledge graphs excel in maintaining logical coherence for multi-step reasoning, where natural language often struggles. By leveraging planning algorithms, they identify optimal reasoning paths, enhancing both interpretability and clarity in handling complex information; 3) Effective handling of linguistic diversity. Natural language often presents challenges with varying terminologies for the same entities or relationships. Knowledge graphs address this by enabling efficient entity and relation resolution through off-the-shelf graph algorithms, streamlining the integration of diverse data sources.
@@ IV-B1 Knowledge Graph Construction from Corpus
This research direction aims to transform unstructured text corpora into structured KGs, enabling efficient and precise information retrieval. It heavily relies on Open Information Extraction techniques, which automate the domain-independent extraction of relational facts from large corpora. These techniques include both traditional OIE approaches [109, 110, 111, 112] and advanced OIE methods based on LLMs [113, 114, 115, 116]. However, constructing effective KGs from domain-specific corpora remains challenging due to variations in ontology across domains. A key objective in this area is to incorporate inductive biases to design graph structures that capture essential domain-specific information while minimizing unnecessary complexity.
@@
IV-B2 GraphRAG with existing KGs
IV-B3 Limitations
Although using graphs as knowledge carriers is efficient and effective, these methods are limited in several aspects: $(i)$ Lack of high-quality KGs. For directly using KGs as external knowledge bases, this line of research is constrained by the availability of high-quality KGs. Constructing KGs is resource-intensive, and although powerful KG refinement techniques [127, 128, 129, 130, 131, 132] have been developed to improve KG completeness, most publicly available KGs remain far from comprehensive. Furthermore, the absence of a unified ontology poses challenges to designing transferable planning algorithms. These limitations have motivated recent research to focus on constructing KGs from corpora. However, this line of work introduces additional challenges: $(ii)$ Trade-off between efficiency and effectiveness. When constructing KGs from text corpora, the granularity of the extracted knowledge plays a crucial role in balancing efficiency and effectiveness. Preserving fine-grained information results in larger, more detailed KGs, which can hinder computational efficiency. Conversely, compact KGs may sacrifice important details, leading to potential information loss. This trade-off complicates the knowledge summarization process. Moreover, approaches like GraphRAG [25] and GraphReader [40], which leverage LLMs for knowledge summarization, face additional challenges due to the high computational costs associated with LLMs.
IV-C Hybrid GraphRAG
V Knowledge Retrieval
Figure 5: Illustration of the overall pipeline of knowledge retrieval.
V-A Overall Pipeline of Knowledge Retriever
The knowledge retrieval process in GraphRAG focuses on extracting relevant background knowledge from a graph database in response to a given query. This process follows three distinct and sequential steps that transform raw graph data into usable, contextual knowledge. The overall process of knowledge retrieval is shown in Figure 5.
V-A1 Query/Graph Preprocessing
V-A2 Matching
V-A3 Knowledge Pruning
V-B Retrieval Techniques
In this subsection, we provide a comprehensive overview of the various retrieval techniques employed in the GraphRAG knowledge base. These techniques can be broadly categorized into six main classes, including string matching, semantic similarity, logical reasoning, LLM-based approaches, RL-based methods, and GNN-based models, as shown in Table II.
V-B1 Semantics Similarity-based Retriever
The semantic similarity-based retriever primarily conducts appropriate answer retrieval by measuring the similarity between queries in discrete linguistic space or continuous vector space and the knowledge base [50, 32, 52, 51, 34]. (i) Discrete space modeling. Discrete space modeling methods primarily leverage linguistic discrete statistical knowledge to directly model text strings, which is a kind of simple and straightforward retrieval technique that relies on the exact matching of query terms with entities and relations in the graph base. It employs algorithms like substring matching, regular expressions, and exact phrase matching to identify relevant nodes and edges [60, 133, 84, 51, 23]. Although straightforward to implement, string matching is limited by its inability to handle variations in terminology, synonyms, and contextual nuances. Despite these limitations, it serves as a foundational approach, especially useful in domains with well-defined vocabularies and minimal ambiguity. (ii) Embedding space modeling. Embedding space modeling techniques are employed to evaluate the contextual and conceptual connections between queries and the elements of a graph base to transcend the constraints of string matching [50, 52, 34, 32]. Utilizing approaches like pre-train language models and word embeddings, such as TF-IDF, Word2Vec, and GloVe [134, 135], these methods capture the semantic ties that bind different components. By mapping both queries and graph entities into a continuous vector space, these sophisticated techniques enable the retrieval of semantically pertinent information, bridging gaps even in the absence of precise term matches. However, the limitation of semantics similarity-based models is that they neglect the full exploitation of graph structures resulting in a significant underutilization of the inherent advantages of graph bases.
TABLE II: Representative knowledge retrieval techniques and strategies used in different GraphRAG Systems.
@@ V-B1 Semantics Similarity-based Retriever
Retrieval Techniques Similarity-based StruGraphRAG [50] query embedding entity embedding BERT/BERT similarity calculation literal context
@@
Retrieval Strategy Multi-round DialogGSR [62] query embedding linearized graph T5/T5 subgraph generation linearized subgraph

### v2

@@ III-A3 Knowledge Integration
The quality of retrieved content can drastically affect generation, with irrelevant or misleading data potentially having deleterious effects. LeanContext [113] addresses this issue by utilizing reinforcement learning to selectively choose sentences that are most pertinent to the query, effectively minimizing the context size and reducing computational costs. Similarly, SELF-RAG [16] introduces a self-reflection mechanism where the language model assesses its own generated and retrieved content, allowing for corrective measures to be implemented during the generation process. SKR [114] proposes a dynamic framework whereby LLMs can rely on their pre-trained knowledge for recognized queries, resorting to retrieval only when necessary. FILCO [115] enhances this by training a context-filtering model that screens out irrelevant data, thereby mitigating hallucination risks. Recent research by Lyu et al. [116] introduces an evaluation method specifically designed to assess the importance of the retrieved content. Their findings suggest that by pruning or re-weighting parts of the retrieval corpus, RAG systems can enhance their performance without requiring additional training rounds. In scenarios where the retrieval corpus is limited, approaches such as Selfmem [117] construct a memory pool from LLM-generated results, using an iterative selection framework to enhance generative quality. SAIL [118] forms an instruction-tuning dataset based on retrieval outcomes, fine-tuning LLMs to ground responses in reliable content while excluding distracting elements.
@@ III-B1 Complex query understanding
Traditional RAG faces significant challenges in precisely answering complex queries, mainly due to the intrinsic limitation of its knowledge organization (vector database). Given a query, these RAG methods only retrieve semantically similar chunks, within which the local contextual information may be insufficient to answer multi-hop questions. This limitation becomes more pronounced as chunk granularity decreases, making it challenging to handle domain knowledge, which often requires multi-hop reasoning for effective understanding. Some traditional RAG methods have attempted to improve the complex query understanding, such as enhancing queries before retrieval [41], enhancing multiple query candidates for re-ranking [42], and using related metadata to provide richer information [35]. However, these methods are still limited by the chunking style of knowledge organization and are challenging to capture multi-hop information in complex queries effectively.
@@ III-C1 Definition of GraphRAG
GraphRAG can be formally defined as a subclass of RAG framework that leverage graph structure to organize and retrieve knowledge. Unlike traditional RAG methods, which rely on vector databases for knowledge organization, GraphRAG employs structural databases where graphs are used to model dependencies among knowledge pieces. This approach enhances the accuracy and efficiency of information retrieval, enabling more effective augmented generation for LLMs. Specifically, GraphRAG systems can utilize graphs as either the carrier of knowledge or the indexing tools for efficient retrieval from chunked textual data. By modeling dependencies between nodes, GraphRAG enables the discovery of related knowledge centered around a topic or anchor entity, ensuring comprehensive knowledge retrieval. Moreover, these connections support efficient search by navigating through relevant pathways and meanwhile pruning irrelevant information during the retrieval process.
@@ III-C2 Workflow of GraphRAG
Knowledge Organization The knowledge organization stage structures external information using graph-based methods, either through explicit knowledge representation (graphs as knowledge carriers) or indexing mechanisms (graphs for knowledge indexing).
Knowledge Retrieval. GraphRAG employs graph-based retrievers to retrieve relevant information based on the input query. This retriever not only considers the semantic similarity between the query and each text chunk but also the logical coherence between the query type and the retrieved subgraph.
Knowledge Integration. Once the relevant knowledge is retrieved, the GraphRAG model integrates it with the input query to generate the final output. The goal of the integration process is to seamlessly incorporate the retrieved knowledge into the generated text, thereby enhancing its quality and informativeness. A key design consideration is how to preserve the richness of the retrieved subgraph information in the final text-based prompt without introducing redundancy or misplacing emphasis on less critical aspects of the context.
@@ III-D Traditional RAG vs GraphRAG
GraphRAG provides several key advantages over traditional RAG systems, enhancing the capabilities of AI-driven information retrieval and generation. Below, we discuss these benefits and summarize the key points in Figure 4.
@@ III-D3 Efficiency and Scalability
GraphRAG systems built on fast graph databases can handle large-scale datasets efficiently. Graph databases are optimized for relationship-based queries, allowing for quick traversal of complex data structures. This efficiency translates to faster response times, especially for queries that require exploring multiple relationships. Research has shown that GraphRAG systems can generate LLM responses using 26% to 97% fewer tokens compared to traditional methods, indicating significant improvements in both speed and resource utilization. GraphRAG systems can more easily accommodate updates to the knowledge base. New information can be added as nodes or edges in the graph without requiring a complete reindexing of the entire knowledge base. This feature allows for real-time updates and ensures that the system always has access to the most current information. The graph structure allows GraphRAG to consider the context of a query more effectively. Instead of relying solely on keyword matching or vector similarity, it can leverage the semantic relationships in the graph to retrieve more relevant information. This context awareness leads to more accurate and pertinent responses.
@@ III-D4 Interpretability
The graph structure in GraphRAG allows for better visualization and understanding of how the system arrives at its responses. Users can trace the path of reasoning through the knowledge graph, seeing which entities and relationships were considered in formulating the answer. This transparency is crucial for building trust in AI systems and is especially valuable in fields like healthcare, finance, or legal applications where decision-making processes need to be auditable.
@@ IV Knowledge Organization
A major distinction between GraphRAG techniques and traditional RAG methods lies in their ability to leverage graph structures for efficient knowledge organization, enhancing the effectiveness of query responses. In this setup, the LLM functions as an intelligent agent, while the graph structures support its ability to organize and integrate information more comprehensively and precisely. In real-world applications, external knowledge sources may include extensive text corpora, document collections, search results, historical user data, or interaction logs. Properly organizing these sources is essential, as direct retrieval can be prone to irrelevant results and missing contexts, leading to information overload, knowledge conflicts, and compromised comprehension.
@@ IV-A Graphs for Knowledge Indexing
The inherent challenges associated with the construction and maintenance of these graphs raise several research questions, giving rise to several research questions that shape current research directions: (1) Conciseness and Relevance: Ensuring the constructed graph captures only relevant relationships without overloading it with unnecessary connections is a significant challenge, thus facilitating effective recalling of relevant text chunks without redundancies, (2) Consistency and Conflict Resolution: Different data chunks may introduce conflicting information. It is crucial to resolve these conflicts and ensure that the graph remains consistent, reliable, and well-structured. Researchers are actively exploring tailored index graph structures to address these challenges effectively.
@@ IV-B Graphs as Knowledge Carriers
The paradigm of representing knowledge explicitly through graph structures is increasingly recognized for its effectiveness, offering several notable advantages: (1) Efficient retrieval of query-related knowledge facts. The explicit structure of knowledge graphs [125] facilitates logic-guided chain retrieval, efficiently identifying missing facts while pruning the search space through reasoning paths; (2) Coherent multi-step reasoning over long spans. Previous studies have revealed the importance of the reasoning steps, i.e., lengthening the reasoning steps in prompts brought significant reasoning performance enhancement [126]. Knowledge graphs excel in maintaining logical coherence for multi-step reasoning, where natural language often struggles. By leveraging planning algorithms, they identify optimal reasoning paths, enhancing both interpretability and clarity in handling complex information; (3) Linguistic diversity. Natural language often presents challenges with varying terminologies for the same entities or relations. KGs address this by enabling efficient entity and relation resolution through off-the-shelf graph algorithms, streamlining the integration of diverse data sources.
@@ IV-B1 Knowledge Graph Construction from Corpus
This research direction aims to transform unstructured text corpora into structured KGs, enabling efficient and precise information retrieval. It heavily relies on Open Information Extraction techniques, which automate the domain-independent extraction of relational facts from large corpora. These techniques include both traditional OIE approaches [127, 128, 129, 130] and advanced OIE methods based on LLMs [131, 132]. However, constructing effective KGs from domain-specific corpora remains challenging due to variations in ontology across domains. A key objective in this area is to incorporate inductive biases to design graph structures that capture essential domain-specific information while minimizing unnecessary complexity.
@@
TABLE I: Representative knowledge retrieval techniques and strategies used in different GraphRAG Systems.
@@
Retrieval Techniques
Similarity-based StruGraphRAG [78] query embedding entity embedding BERT/BERT similarity calculation literal context
@@
Retrieval Strategy
Multi-round DialogGSR [92] query embedding linearized graph T5/T5 subgraph generation linearized subgraph
@@
IV-B2 GraphRAG with existing KGs
IV-B3 Limitations
Although using graphs as knowledge carriers is efficient and effective, these methods are limited in several aspects: $(i)$ Lack of high-quality KGs. For directly using KGs as external knowledge bases, this line of research is constrained by the availability of high-quality KGs. Constructing KGs is resource-intensive, and although powerful KG refinement techniques [143, 144, 145, 146, 147, 148] have been developed to improve KG completeness, most publicly available KGs remain far from comprehensive. Furthermore, the absence of a unified ontology poses challenges to designing transferable planning algorithms. These limitations have motivated recent research to focus on constructing KGs from corpora. However, this line of work introduces additional challenges: $(ii)$ Trade-off between efficiency and effectiveness. When constructing KGs from text corpora, the granularity of the extracted knowledge plays a crucial role in balancing efficiency and effectiveness. Preserving fine-grained information results in larger-scale KGs, which may hinder computational efficiency. Conversely, compact KGs may sacrifice important details, leading to potential information loss. This trade-off complicates the knowledge summarization process. Moreover, approaches like GraphRAG [24], which leverage LLMs for knowledge summarization, suffer from the high token costs and unbearable high runtime overhead.
IV-C Hybrid GraphRAG
V Knowledge Retrieval
V-A Overall Pipeline of Knowledge Retriever
The knowledge retrieval process in GraphRAG focuses on extracting relevant background knowledge from a graph database in response to a given query. This process follows three distinct and sequential steps that transform raw graph data into usable, contextual knowledge.
V-A1 Query/Graph Preprocessing
V-A2 Matching
V-A3 Knowledge Pruning
V-B Retrieval Techniques
In this subsection, we provide a comprehensive overview of the various retrieval techniques employed in the GraphRAG knowledge base. These techniques can be broadly categorized into six main classes, including string matching, semantic similarity, logical reasoning, LLM-based approaches, RL-based methods, and GNN-based models, as shown in Table I.
V-B1 Semantics Similarity-based Retriever
The semantic similarity-based retriever primarily conducts appropriate answer retrieval by measuring the similarity between queries in discrete linguistic space or continuous vector space and the knowledge base [78, 45, 80, 79, 47]. (i) Discrete space modeling. Discrete space modeling methods primarily leverage linguistic discrete statistical knowledge to directly model text strings, which is a kind of simple and straightforward retrieval technique that relies on the exact matching of query terms with entities and relations in the graph base. It employs algorithms like substring matching, regular expressions, and exact phrase matching to identify relevant nodes and edges [88, 34, 79, 22]. Although straightforward to implement, string matching is limited by its inability to handle variations in terminology, synonyms, and contextual nuances. Despite these limitations, it serves as a foundational approach, especially useful in domains with well-defined vocabularies and minimal ambiguity. (ii) Embedding space modeling. Embedding space modeling techniques are employed to evaluate the contextual and conceptual connections between queries and the elements of a graph base to transcend the constraints of string matching [78, 80, 47, 45]. Utilizing approaches like pre-train language models and word embeddings, such as TF-IDF, Word2Vec, and GloVe [149, 150], these methods capture the semantic ties that bind different components. By mapping both queries and graph entities into a continuous vector space, these sophisticated techniques enable the retrieval of semantically pertinent information, bridging gaps even in the absence of precise term matches. However, the limitation of semantics similarity-based models is that they neglect the full exploitation of graph structures resulting in a significant underutilization of the inherent advantages of graph bases.